OpenStudy (anonymous):

Simplify: (2x^2 y^4)^3 -8x5y7 -6x6y12 none of the above -6x5y7

OpenStudy (shamil98):

Use the rule \[\huge (x^m)^n = x^{n*m}\]

OpenStudy (anonymous):

Simplify: (2x^2 y^4)^3 A.-8x5y7 B.-6x6y12 C.none of the above D.-6x5y7

OpenStudy (shamil98):

Take everything in the parentheses to the power of 3.

OpenStudy (shamil98):

\[\huge 2^3 x^{2*3}y^{4*3}\]

OpenStudy (anonymous):

Use this: \[\left( x^{2}*y^{3} \right)^{k} = \left( x^{2*k}*y^{3*k} \right)\]

OpenStudy (anonymous):

so its none of these?

OpenStudy (anonymous):

ok, simplify shamil98's response by multiplying the exponents and finding the 2^3